theres a jar of nickels and dimes that has $6.75. theres 84 more nickels than dimes. how many nickels and how many dimes are there?
Answer by checkley71(8403)

5N+10D=675
N=D+84 NOW SUBSTITUTE(D+84) FOR N IN THE FIRST EQUATION THUS
5(D+84)+10D=675
15D+420=675
15D=675-420
15D=255
D=255/15
D=17 NUMBER OF DIMES THEN THE NICKELS ARE 17+84=101
PROOF
5*101+17*10=675
505+170=675
675=675